# Privacy-preserving Blockchain-enabled Parametric Insurance via Remote Sensing and IoT

### v2

@@ Abstract
Traditional Insurance, a popular approach of financial risk management, has suffered from the issues of high operational costs, opaqueness, inefficiency and a lack of trust. Recently, blockchain-enabled parametric insurance through authorized data sources (e.g., remote sensing and IoT) aims to overcome these issues by automating the underwriting and claim processes of insurance policies on a blockchain. However, the openness of blockchain platforms raises a concern of user privacy, as the private user data in insurance claims on a blockchain may be exposed to outsiders. In this paper, we propose a privacy-preserving parametric insurance framework based on succinct zero-knowledge proofs (zk-SNARKs), whereby an insuree submits a zero-knowledge proof (without revealing any private data) for the validity of an insurance claim and the authenticity of its data sources to a blockchain for transparent verification. Moreover, we extend the recent zk-SNARKs to support robust privacy protection for multiple heterogeneous data sources and improve its efficiency to cut the incurred gas cost by 80%. As a proof-of-concept, we implemented a working prototype of bushfire parametric insurance on real-world blockchain platform Ethereum, and present extensive empirical evaluations. [^1] This is an extended version of the journal paper to appear in IEEE Trans. Services Computing [1].
@@ I Introduction
Insurance is a financial agreement that provides protection against potential financial losses resulting from specific future risks. Essentially, an insurer agrees to compensate an insuree (i.e., client) for potential losses resulting from specific future events, in exchange for a premium paid by the insuree. This enables the transfer of uncertain financial loss risks from an insuree to an insurer. Nowadays, there are a wide range of insurance products, covering natural events, health care, and diverse personal or commercial circumstances.
@@ I-B Blockchain-enabled Parametric Insurance
A smart contract is an executable script that runs atop a blockchain without a third party, and can be triggered once a set of predetermined conditions are met. Compared to traditional paper-based contracts, the execution of smart contracts is publicly verifiable and non-retractable, reinforcing transparency and accountability in a trustless environment. These properties are ideal for realizing parametric insurance.
@@
Note that IoT sensors and remote sensing can capture different data from the environment to support parametric insurance. For example, IoT sensors can measure local wind speed, rainfall, temperature and air quality indexes, while remote sensing can identify spatial events like bushfires and droughts. There are a number of publicly available sensor data repositories available for public access (e.g., from NASA [6] or ESA [7] satellites), allowing an insuree to conduct self-verification of parametric insurance claims.
@@ II-A Blockchain-based Insurance
  3. 3.
    Off-chain Databases: There are off-chain databases hosted by a third-party, which carries out off-chain data validation such that only the final decisions of claims are stored on blockchain without revealing insurees’ privacy. Nevertheless, this approach lacks transparency and accountability since claim validation is not conducted on blockchain. As with traditional insurance, insurees must place their trust in the third-party’s off-chain operations.
To the best of our knowledge, there is no other existing decentralized insurance solution that provides strong privacy protection, while allowing transparent on-chain verification of insurance claims. Our solution is the first study to provide a practical privacy-preserving and publicly verifiable solution that is implemented on real-world permissionless blockchain.
@@ III Cryptographic Preliminaries
First, ${\mathbb{F}}_{p}=\{0,...,p-1\}$ denote a finite field of integers modulo $p$. We write “$x+y$” and “$xy$” for modular arithmetic without explicitly mentioning “${\tt mod\ }p$”. We consider a cyclic group ${\mathbb{G}}$ of prime order $p$ (e.g., an elliptic curve group). Let ${\tt g}$ be a generator of ${\mathbb{G}}$, such that ${\tt g}$ can generate any element in ${\mathbb{G}}$ by taking proper powers (i.e., for each ${\tt k}\in{\mathbb{G}}$, there exists $x\in{\mathbb{F}}_{p}$ such that ${\tt k}={\tt g}^{x}$). We write $x\xleftarrow{\$}\mathbb{\mathbb{F}}_{p}$ to mean selecting $x$ in ${\mathbb{F}}_{p}$ at uniformly random. The computational Diffie-Hellman assumption states that given ${\tt g}^{x}$, it is computationally hard to obtain $x$, which underlies the security of many crypto systems.
@@ III-B Polynomial Commitment
  • Setup$({\lambda})\rightarrow{\tt srs}$: Setup takes the security parameter ${\lambda}$ and outputs a structured reference string ${\tt srs}$, which is a public parameter to the commitment scheme.
  • Commit$({\tt srs},f[X])\rightarrow F$: Commit generates a commitment $F$ given a polynomial $f[X]$.
  • Open$({\tt srs},F,z)\rightarrow(v,\pi)$: Open evaluates polynomial $f[X]$ with its commitment $F$ at value $X=z$, and outputs evaluation $v=f[z]$ (i.e., an opening), together with a proof $\pi$ to prove the following relation:
@@
  • Verify$({\tt srs},F,z,v,\pi)\rightarrow\{{\tt True},{\tt False}\}$: Verify uses proof $\pi$ to verify whether $(F,z,v)\in{\mathscr{R}}_{\rm cm}$. It outputs ${\tt True}$, if the verification is passed, otherwise ${\tt False}$.
@@ III-C KZG Polynomial Commitment
Given a Laurent polynomial $f[X]$ and value $z$, then the polynomial $f[X]-f[z]$ is divisible by $X-z$, namely, $f[X]-f[z]=(X-z)\cdot q[X]$ for some Laurent polynomial $q[X]$. Intuitively, it is because $X=z$ is a root of $f[X]-f[z]$.
@@ III-D Digital Signature
  • Setup$({\lambda})\rightarrow({\tt pk,sk})$: Setup takes a security parameter ${\lambda}$ and outputs an asymmetric key pair $({\tt pk,sk})$.
  • Sign$({\tt sk},m)\rightarrow\sigma$: Sign takes a message $m$ with a fixed length, and outputs a signature $\sigma$ using the secret key ${\tt sk}$.
  • VerifySign$({\tt pk},m,\sigma)\rightarrow\{{\tt True},{\tt False}\}$: VerifySign uses the public key ${\tt pk}$ to verify message $m$ and signature $\sigma$. It outputs ${\tt True}$, if the verification is passed, otherwise ${\tt False}$.
@@ IV-A Arguments of Knowledge and zk-SNARKs
  • Setup$(\lambda,C)\rightarrow{\tt srs}$: Setup takes a relation ${\mathscr{R}}_{C}$ as well as a security parameter $\lambda$ as input, and outputs a structured reference string containing the public parameters for proof generation and verification.
  • Prove$({\tt srs},x,w)\rightarrow\pi$: Prove uses the common reference string ${\tt srs}$, public input $x$, and the secret witness $w$ to generate a proof $\pi$, which is a proof of the following statement: “given $C$ and $x$, there exists a secret witness $w$, such that $(x,w)\in{\mathscr{R}}_{C}$”.
  • Verify$({\tt srs},x,\pi)\rightarrow\{{\tt True},{\tt False}\}$: Verify verifies the proof $\pi$, and outputs ${\tt True}$ if $(x,w)\in{\mathscr{R}}_{C}$, otherwise ${\tt False}$.
@@
A generic argument system consists of three methods (Setup, Prove, Verify), as explained in Table IV. An argument system has several key properties, typically, completeness, knowledge soundness and perfect honest-verifier zero knowledge. Zero-Knowledge Succinct Non-Interactive Arguments of Knowledge (zk-SNARKs) are argument systems with the additional properties of succinctness and non-interactiveness. We refer the readers to the Appendix for the formal definitions of these properties.
@@ IV-C1 Mapping Decision Function to Constraint System
$\mathbf{a}_{i}\cdot\mathbf{b}_{i}&=\mathbf{c}_{i},&\mbox{for\ }1\leq i\leq N\\ \mathbf{a}\cdot\mathbf{u}_{q}+\mathbf{b}\cdot\mathbf{v}_{q}+\mathbf{c}\cdot\mathbf{w}_{q}&=k_{q},&\mbox{for\ }1\leq q\leq Q\\ \end{array}\right.\qquad(\mathscr{C})$
@@ IV-C2 Mapping Constraint System to Polynomial IOP
$\hat{\mathbf{u}}_{i}[Y]\triangleq$ $\sum^{Q}_{q=1}\mathbf{u}_{q,i}Y^{q+N},\qquad\qquad\qquad\hat{\mathbf{v}}_{i}[Y]\triangleq\sum^{Q}_{q=1}\mathbf{v}_{q,i}Y^{q+N}$
$\hat{\mathbf{w}}_{i}[Y]\triangleq$ $-Y^{i}-Y^{-i}+\sum^{Q}_{q=1}\mathbf{w}_{q,i}Y^{q+N},\ \hat{k}[Y]\triangleq\sum^{Q}_{q=1}k_{q}Y^{q+N}$
@@
$r[X,Y]\triangleq$ $\sum^{N}_{i=1}{\mathbf{a}}_{i}(XY)^{i}+\sum^{N}_{i=1}{\mathbf{b}}_{i}(XY)^{-i}+\sum^{N}_{i=1}{\mathbf{c}}_{i}(XY)^{-i-N}$
$s[X,Y]\triangleq$ $\sum^{N}_{i=1}\hat{\mathbf{u}}_{i}[Y]X^{-i}+\sum^{N}_{i=1}\hat{\mathbf{v}}_{i}[Y]X^{i}+\sum^{N}_{i=1}\hat{\mathbf{w}}_{i}[Y]X^{i+N}$
$t[X,Y]\triangleq$ $r[X,1]\big{(}r[X,Y]+s[X,Y]\big{)}-\hat{k}[Y]$ (4)
@@ IV-C3 Sonic Protocol
    Additional Blinders: Because there are three values revealed related to $r[x,y]$, which includes $r[z,1]$, $r[zy,1]$, and $\textsf{Commit}({\tt srs},r[X,1])$, we add four blinders to the polynomial with random coefficients and powers ($-2n-1,-2n-2,-2n-3,-2n-4$). Such blinders make the polynomial indistinguishable from any random polynomial given less than four evaluations. Since the random blinders are also part of user input (like other coefficients of $r$), it does not affect the rest of the protocol.
@@
    Checking Zero Constant Term: Note that we cannot check if $t[X,y]$ has a zero constant term by checking $t[0,y]\stackrel{{\scriptstyle?}}{{=}}0$, because $t[X,y]$ is a Laurent polynomial, which is undefined at $t[0,y]$. To resolve this issue, we use a restricted version of KZG polynomial commitment scheme (see Appendix VIII-A) that precludes a committed polynomial with a non-zero constant term, and hence, forcing the prover to only commit polynomials with zero constant terms. Step (5) in the above protocol can be skipped.
@@
    $(r_{1},\pi_{r_{1}})&\leftarrow{\textsf{\scriptsize Open}}({\tt srs},R,z),&(r_{2},\pi_{r_{2}})&\leftarrow{\textsf{\scriptsize Open}}({\tt srs},R,zy)\\ (t,\pi_{t})&\leftarrow{\textsf{\scriptsize Open}}({\tt srs},T,z)\end{array}$
@@
    {$(k,\pi_{k})&\leftarrow{\textsf{\scriptsize Open}}({\tt srs},K,y),&(s,\pi_{s})&\leftarrow{\textsf{\scriptsize Open}}({\tt srs},S_{X},z)\\ (s_{1},\pi_{s_{1}})&\leftarrow{\textsf{\scriptsize Open}}({\tt srs},S_{X},1),&(s_{2},\pi_{s_{2}})&\leftarrow{\textsf{\scriptsize Open}}({\tt srs},S_{Y},y)\\ &t\leftarrow{r_{1}(r_{2}+s)-k}\end{array}\right.$
@@ V-A Authenticating Heterogeneous Data Sources
We denote the intermediate data in the computation by $(\tilde{\mathbf{a}},\tilde{\mathbf{b}},\tilde{\mathbf{c}})$. Also, we denote the polynomial for encoding $(\tilde{\mathbf{a}},\tilde{\mathbf{b}},\tilde{\mathbf{c}})$ by $\tilde{r}[X,Y]$, which also includes the random blinders as described in Section IV-C3. However, because there are in total $5+2j$ evaluations of $r[X,Y]$ revealed ($2+j$ commitments and $3+j$ openings), we set $6+2j$ random blinders to preserve the security of the protocol:
$\tilde{r}[X,Y]=\tilde{r}[X,Y]+\sum^{6+2j}_{i=1}\tilde{\mathbf{g}}_{i}(XY)^{-i-2N}\$
@@
The rest of the protocol is similar to the one in Table V. We leave the security proof of the protocol in the Appendix.
@@ V-B Polynomial Commitments with Batch Verification
Sonic protocol requires the openings of multiple polynomials at different evaluation points at the verification stage. Normally, this is implemented by separate verification operations. However, it is possible to reduce the verification overhead by batch verification of multiple openings together. We propose a new polynomial commitment scheme specifically designed for Sonic to allow simultaneous openings of multiple polynomials at different evaluation points. Formally, we consider a set of $K$ polynomials $\{f_{i}[X]\}_{i=1}^{K}$. The prover commits to these polynomials as $F_{i}=\textsf{\footnotesize Commit}({\tt srs},f_{i}[X])$. There are a set of evaluation points ${\mathcal{S}}=\{z_{1},...,z_{n}\}$. Only a subset of evaluation points ${\mathcal{S}}_{i}\subset{\mathcal{S}}$ are evaluated on the $i$-th polynomial $f_{i}[X]$. The prover can open $\{(f_{i}[z])_{z\in{\mathcal{S}}_{i}}\}_{i=1}^{K}$ in a batch to the verifier with one single proof to prove all the openings are correct with respect to the commitments $(F_{i})_{i=1}^{K}$. We defer the details to the Appendix. In Appendix VIII-B, we will describe the new polynomial commitment scheme with a specific requirement in Sonic protocol to preclude a committed polynomial with a non-zero constant term. In Appendix VIII-D, we will describe an enhanced Sonic protocol with batch verification (${\tt S}_{\tt EV}$).
@@
    Verifier $\Rightarrow$ Prover: $y\xleftarrow{\$}\mathbb{\mathbb{F}}_{p}$ // (Fiat-Shamir): $y\leftarrow{\textsf{\scriptsize Hash}}(R|\tilde{R})$
@@
    {$(k,\pi_{k})&\leftarrow{\textsf{\scriptsize Open}}({\tt srs},K,y),&(\hat{s},\pi_{\hat{s}})&\leftarrow{\textsf{\scriptsize Open}}({\tt srs},S_{X},z)\\ (\hat{s}_{1},\pi_{\hat{s}_{1}})&\leftarrow{\textsf{\scriptsize Open}}({\tt srs},S_{X},1),&(\hat{s}_{2},\pi_{\hat{s}_{2}})&\leftarrow{\textsf{\scriptsize Open}}({\tt srs},S_{Y},y)\end{array}\right.$

### v3

@@ Abstract
Traditional Insurance, a popular approach of financial risk management, has suffered from the issues of high operational costs, opaqueness, inefficiency and a lack of trust. Recently, blockchain-enabled parametric insurance through authorized data sources (e.g., remote sensing and IoT) aims to overcome these issues by automating the underwriting and claim processes of insurance policies on a blockchain. However, the openness of blockchain platforms raises a concern of user privacy, as the private user data in insurance claims on a blockchain may be exposed to outsiders. In this paper, we propose a privacy-preserving parametric insurance framework based on succinct zero-knowledge proofs (zk-SNARKs), whereby an insuree submits a zero-knowledge proof (without revealing any private data) for the validity of an insurance claim and the authenticity of its data sources to a blockchain for transparent verification. Moreover, we extend the recent zk-SNARKs to support robust privacy protection for multiple heterogeneous data sources and improve its efficiency to cut the incurred gas cost by 80%. As a proof-of-concept, we implemented a working prototype of bushfire parametric insurance on real-world blockchain platform Ethereum, and present extensive empirical evaluations.
@@ II-A Blockchain-based Insurance
To the best of our knowledge, there is no other existing decentralized insurance solution that provides strong privacy protection, while allowing transparent on-chain verification of insurance claims.
@@ III Cryptographic Preliminaries
First, ${\mathbb{F}}_{p}=\{0,...,p-1\}$ denote a finite field of integers modulo $p$. We write “x+yx+y” and “xyxy” for modular arithmetic without explicitly mentioning ${\tt mod\ }p$”. We consider a cyclic group $\mathbb{G}$} of prime order $p$ (e.g., an elliptic curve group). Let $\tt g$} be a generator of $\mathbb{G}$}, such that $\tt g$} can generate any element in $\mathbb{G}$} by taking proper powers (i.e., for each ${\tt k}\in{\mathbb{G}}$, there exists $x\in{\mathbb{F}}_{p}$ such that ${\tt k}={\tt g}^{x}$). We write $x\xleftarrow{\$}\mathbb{\mathbb{F}}_{p}$ to mean selecting $x$ in ${\mathbb{F}}_{p}$ at uniformly random. The computational Diffie-Hellman assumption states that given ${\tt g}^{x}$, it is computationally hard to obtain $x$, which underlies the security of many crypto systems.
@@ III-B Polynomial Commitment
  • $({\lambda})\rightarrow{\tt srs}: Setup$ takes the security parameter $\lambda$} and outputs a structured reference string $\tt srs$}, which is a public parameter to the commitment scheme.
  • $({\tt srs},f[X])\rightarrow F: Commit$ generates a commitment $F$ given a polynomial $f[X]$.
  • $({\tt srs},F,z)\rightarrow(v,\pi): Open$ evaluates polynomial $f[X]$ with its commitment $F$ at value $X=z$, and outputs evaluation $v=f[z]$ (i.e., an opening), together with a proof $\pi$ to prove the following relation:
@@
  • $({\tt srs},F,z,v,\pi)\rightarrow\{{\tt True},{\tt False}\}: Verify$ uses proof $\pi$ to verify whether $(F,z,v)\in{\mathscr{R}}_{\rm cm}$. It outputs $\tt True$}, if the verification is passed, otherwise $\tt False$}.
@@ III-C KZG Polynomial Commitment
Given a Laurent polynomial $f[X]$ and value $z$, then the polynomial f[X]f[z]f[X]-f[z] is divisible by XzX-z, namely, $f[X]-f[z]=(X-z)\cdot q[X]$ for some Laurent polynomial $q[X]$. Intuitively, it is because $X=z$ is a root of f[X]f[z]f[X]-f[z].
@@ III-D Digital Signature
  • $({\lambda})\rightarrow({\tt pk,sk}): Setup$ takes a security parameter $\lambda$} and outputs an asymmetric key pair $({\tt pk,sk})$.
  • $({\tt sk},m)\rightarrow\sigma: Sign$ takes a message $m$ with a fixed length, and outputs a signature $\sigma$ using the secret key $\tt sk$}.
  • $({\tt pk},m,\sigma)\rightarrow\{{\tt True},{\tt False}\}: VerifySign$ uses the public key $\tt pk$} to verify message $m$ and signature $\sigma$. It outputs $\tt True$}, if the verification is passed, otherwise $\tt False$}.
@@ IV-A Arguments of Knowledge and zk-SNARKs
  • $(\lambda,C)\rightarrow{\tt srs}: Setup$ takes a relation ${\mathscr{R}}_{C}$ as well as a security parameter $\lambda$ as input, and outputs a structured reference string containing the public parameters for proof generation and verification.
  • $({\tt srs},x,w)\rightarrow\pi: Prove$ uses the common reference string $\tt srs$}, public input $x$, and the secret witness $w$ to generate a proof $\pi$, which is a proof of the following statement: “given $C$ and $x$, there exists a secret witness $w$, such that $(x,w)\in{\mathscr{R}}_{C}$”.
  • $({\tt srs},x,\pi)\rightarrow\{{\tt True},{\tt False}\}: Verify$ verifies the proof $\pi$, and outputs $\tt True$} if $(x,w)\in{\mathscr{R}}_{C}$, otherwise $\tt False$}.
@@
A generic argument system consists of three methods (Setup, Prove, Verify), as explained in Table IV. An argument system has several key properties, typically, completeness, knowledge soundness and perfect honest-verifier zero knowledge. Zero-Knowledge Succinct Non-Interactive Arguments of Knowledge (zk-SNARKs) are argument systems with the additional properties of succinctness and non-interactiveness.
@@ IV-C1 Mapping Decision Function to Constraint System
[…]
@@ IV-C2 Mapping Constraint System to Polynomial IOP
$\hat{\mathbf{u}}_{i}[Y]\triangleq\$$\sum^{Q}_{q=1}\mathbf{u}_{q,i}Y^{q+N},\qquad\qquad\qquad\hat{\mathbf{v}}_{i}[Y]\triangleq\sum^{Q}_{q=1}\mathbf{v}_{q,i}Y^{q+N}$
$\hat{\mathbf{w}}_{i}[Y]\triangleq\$$-Y^{i}-Y^{-i}+\sum^{Q}_{q=1}\mathbf{w}_{q,i}Y^{q+N},\ \hat{k}[Y]\triangleq\sum^{Q}_{q=1}k_{q}Y^{q+N}$
@@
$r[X,Y]\triangleq\$$\sum^{N}_{i=1}{\mathbf{a}}_{i}(XY)^{i}+\sum^{N}_{i=1}{\mathbf{b}}_{i}(XY)^{-i}+\sum^{N}_{i=1}{\mathbf{c}}_{i}(XY)^{-i-N}$
$s[X,Y]\triangleq\$$\sum^{N}_{i=1}\hat{\mathbf{u}}_{i}[Y]X^{-i}+\sum^{N}_{i=1}\hat{\mathbf{v}}_{i}[Y]X^{i}+\sum^{N}_{i=1}\hat{\mathbf{w}}_{i}[Y]X^{i+N}$
$t[X,Y]\triangleq\$$r[X,1]\big{(}r[X,Y]+s[X,Y]\big{)}-\hat{k}[Y]$ (4)
@@ IV-C3 Sonic Protocol
    Additional Blinders: Because there are three values revealed related to $r[x,y]$, which includes $r[z,1]$, $r[zy,1]$, and $\textsf{Commit}({\tt srs},r[X,1])$, we add four blinders to the polynomial with random coefficients and powers (2n1,2n2,2n3,2n4-2n-1,-2n-2,-2n-3,-2n-4). Such blinders make the polynomial indistinguishable from any random polynomial given less than four evaluations. Since the random blinders are also part of user input (like other coefficients of $r$), it does not affect the rest of the protocol.
@@
    Checking Zero Constant Term: Note that we cannot check if $t[X,y]$ has a zero constant term by checking $t[0,y]\stackrel{{\scriptstyle?}}{{=}}0$, because $t[X,y]$ is a Laurent polynomial, which is undefined at $t[0,y]$. To resolve this issue, we use a restricted version of KZG polynomial commitment scheme (see Appendix A) that precludes a committed polynomial with a non-zero constant term, and hence, forcing the prover to only commit polynomials with zero constant terms. Step (5) in the above protocol can be skipped.
@@
[…]
@@
[…]
@@ V-A Authenticating Heterogeneous Data Sources
[…]
$\tilde{r}[X,Y]=\tilde{r}[X,Y]+\sum^{6+2j}_{i=1}\tilde{\mathbf{g}}_{i}(XY)^{-i-2N}\\$
@@
The rest of the protocol is similar to the one in Table V. We leave the security proof of the protocol in [29].
@@ V-B Polynomial Commitments with Batch Verification
Sonic protocol requires the openings of multiple polynomials at different evaluation points at the verification stage. Normally, this is implemented by separate verification operations. However, it is possible to reduce the verification overhead by batch verification of multiple openings together. We propose a new polynomial commitment scheme specifically designed for Sonic to allow simultaneous openings of multiple polynomials at different evaluation points. Formally, we consider a set of $K$ polynomials $\{f_{i}[X]\}_{i=1}^{K}$. The prover commits to these polynomials as $F_{i}=\textsf{\footnotesize Commit}({\tt srs},f_{i}[X])$. There are a set of evaluation points ${\mathcal{S}}=\{z_{1},...,z_{n}\}$. Only a subset of evaluation points ${\mathcal{S}}_{i}\subset{\mathcal{S}}$ are evaluated on the $i$-th polynomial $f_{i}[X]$. The prover can open $\{(f_{i}[z])_{z\in{\mathcal{S}}_{i}}\}_{i=1}^{K}$ in a batch to the verifier with one single proof to prove all the openings are correct with respect to the commitments $(F_{i})_{i=1}^{K}$. In Appendix B, we describe a new polynomial commitment scheme with a specific requirement in Sonic to preclude a committed polynomial with a non-zero constant term. In [29], we also describe an enhanced Sonic protocol with batch verification (${\tt S}_{\tt EV}$).
@@
[…]
@@
[…]